Graph 4x+18=-2x and define slope, x&y intercept.
I found: 4x=-2x-18=
4x+2x=-2x-18+2x
6x/6=-18/6
x=-3
But slope and y is 0, is this correct? How would I graph this? Thank you

Answer by Alan3354(69443)

Graph 4x+18=-2x and define slope, x&y intercept.
I found: 4x=-2x-18=
4x+2x=-2x-18+2x
6x/6=-18/6
x=-3
But slope and y is 0, is this correct? How would I graph this?
--------------
The slope is not zero.
y can be zero, or any value.
--------------
x = -3 for all values of y.
To graph it, pick any values for y.
eg
(-3,0) and (-3,10)
Draw a line thru the 2 points.
-----
The line is parallel to the y-axis.
The slope is infinite, or undefined.
The x-intercept is -3, or (-3,0)
There is no y-intercept, it's parallel to the y-axis.